A hicker begins a trip by first walking 25 km south east from her car then she walks 40 km in a direction 60 degree centigrade north of East (a) determine the component for each displacement (b) determine R esultant displacement
1
Expert's answer
2020-02-13T08:53:48-0500

a)


"d_{1x}=25\cos{45}=17.68\ km"

"d_{1y}=-25\sin{45}=-17.68\ km"

"d_{2x}=40\cos{60}=20\ km"

"d_{2y}=40\sin{60}=34.64\ km"

b)


"R_x=17.68+20=37.68\ km"

"R_y=-17.68+34.64=16.96\ km"

The magnitude:


"R=\sqrt{37.68^2+16.96^2}=41.3\ km"

The direction:


"\theta=\arctan{\frac{16.96}{37.68}}=24.2\degree N\ of\ E"